F(x) = 5x2 -8x + 10


1
Expert's answer
2022-03-09T07:04:00-0500

"D=b^2-4ac=64-4*50" <0

"\forall x\in\R" F(x)>0

F'(x)=10x-8

F'(x)=0;x=0.8

min(F(x))=F(0.8)="\frac{34}{5}"